# Non-Markovian Quantum Jump Method for Driven-Dissipative Two-Level Systems

### v1

@@ V Summary
The existence probability directly gives the portion of the corresponding quantum trajectory in all the realizations in the limit of infinite $N$. Moreover the normalization of the existence probability is naturally satisfied by definition when all the trajectories are considered. It is shown that the existence probability can be calculated in a top-bottom manner, so in a realistic simulation one can only keep the trajectory class up to a truncation number $n^{*}$ provided that the contributions of those quantum trajectories undergo more than $n^{*}$ normal jumps are negligible to the sum. This makes our method more feasible from the perspective of costing less computational resources.
@@
In future work, it will be interesting to investigate the effect of non-Markovianity on the properties of quantum many-body systems, such as the information scrambling li_PRA_2020 and steady state of spin chains with boundary dissipation popkov_2020 . Finally, we hope the presented method may provide an promising way for efficiently simulating the non-Markovian dynamics of open quantum system.

### v2

@@ VI Summary
The existence probability directly gives the portion of the corresponding quantum trajectory in all the realizations in the limit of infinite $N$. Moreover the normalization of the existence probability is naturally satisfied by definition when all the trajectories are considered. It is shown that the existence probability can be calculated in a top-bottom manner, so in a realistic simulation one can only keep the trajectory class up to a truncation number $n^{*}$ provided that the contributions of those quantum trajectories undergo more than $n^{*}$ normal jumps are negligible to the sum. This makes our method more feasible from the perspective of costing less computational resources. The truncation can usually be adopted in the cases that, for examples, the driving field on the single-body system or the interaction between subsystem of a many-body system are not strong. Because the probability of normal quantum jump is proportional to the population of the upper level of the system and the amplitude of the decay rate. A normal quantum jump projects the system to the lower-level state, thus the population of upper level remains small and the next quantum jump is suppressed. Moreover, at the end of the period of positive decay rate, the amplitude of decay rate approaches to zero which further suppress the probability of normal quantum jump. So a quantum trajectory with multiply quantum jumps is unlikely to appear.
When the truncation is exact, that is there are no trajectories with more than $n^{*}$ jumps, our method coincides with the standard method. However, even if in this situation, our method still has advantages because it gives the explicit expression of the existence probability, while a sufficient large number of realizations are needed in the standard quantum jump method. On the other hand, if the existence probability of the trajectory with more than $n^{*}$ jumps is small but finite, the ensemble average can no longer fully reproduce the density matrix in the Lindblad master equation because part of the information is omitted due to the truncation. Here, we would like to emphasize that the efficient of the modified NMQJ for a more general system is restricted. For instance, in the quantum transport scenario where the quantum system interacting with two unbalanced baths, it is difficult to keep the history of the system because there will be an enormous number of quantum trajectories due to many quantum jumps may occur.
@@
In future work, it will be interesting to investigate the effect of non-Markovianity on the properties of quantum many-body systems, such as the information scrambling li_PRA_2020 and steady state of spin chains with boundary dissipation popkov_2020 . Finally, we hope the presented method may provide a potential way for efficiently simulating the non-Markovian dynamics of open quantum system.